OpenStudy (abbycross167):

20. Which rational number lies between √10 and √12 ? a. 45/14 b. 47/13 c. 47/19 d. 52/17 My Answer ~A 21. A shelter has enough food to feed an average of 80 hungry people a day for 14 days. If an average of 32 people are at the shelter each day, how long will the food last? a. 5.6 days b. 35 days c. 183 days d. 200 days My answer~B 22. Which of the following numbers is a solution of the inequality 9 – x > 2? (NOTE: the > sign has a line under it) a. 7 b. 8 c. 9 d. 10 My Answer~A

OpenStudy (dayakar):

20) \[\sqrt{10}= 3.1\] \[\sqrt{12}=3.4\] approxmate values 45/14 =3.2,, ans:a

OpenStudy (dayakar):

21) a

OpenStudy (dayakar):

22)a